# Performance Comparisons of Reinforcement Learning Algorithms for Sequential Experimental Design

### v1

@@ 3.2 Hidden-Parameter Markov Decision Process
[…]
@@
    where $C_{t}\cdot\boldsymbol{1}$ is a dot product with a vector of “$1$”s;
@@
To find the optimal policy we utilise deep RL algorithms [^3] In the experiments we conduct using these algorithms, we assume $\gamma$ values close to $1$ (rather than $\gamma=1$) for computational stability; this follows Blau et al. (2022). (François-Lavet et al. 2018). The policy is implemented as a deep neural network that relies on a summary of $h_{t}$ in the form of an encoder network. Although the definition of the HiP-MDP satisfies the Markov property, it does so by explicitly including $h_{t}$ as part of $s_{t}$; the use of an encoder network weakens this by not explicitly requiring $h_{t}$ to be part of the state. Instead, the encoder network is recursively defined, so that a summary representation of $h_{t}$ at time $t$ only depends on the summary representation at time $t-1$ and $(\xi_{t},y_{t})$ (Blau et al. 2022). A detailed description of the policy network, as well as more details on the HiP-MDP implementation, are given in Appendix A.
@@ 4 Algorithms
REDQ $6.279\pm 0.013$ $11.689\pm 0.012$ $11.881\pm 0.013$ $11.507\pm 0.015$ $11.095\pm 0.017$ 13.23h
SBR $6.165\pm 0.012$ $11.362\pm 0.013$ $11.788\pm 0.014$ $11.518\pm 0.016$ $11.257\pm 0.017$ 13.52h
DroQ $6.368\pm 0.013$ $11.680\pm 0.013$ $11.902\pm 0.013$ $11.480\pm 0.015$ $11.090\pm 0.017$ 19.05h
SUNRISE $6.340\pm 0.013$ 11.837 $\pm$ 0.012 $12.133\pm 0.013$ 11.846 $\pm$ 0.014 $11.445\pm 0.016$ 21.44h
SUNRISE-DroQ 6.433 $\pm$ 0.012 $11.770\pm 0.012$ 12.143 $\pm$ 0.013 $11.831\pm 0.014$ 11.453 $\pm$ 0.016 29.77h
Random $5.057\pm 0.010$ $7.443\pm 0.016$ $8.820\pm 0.018$ $9.730\pm 0.019$ $10.346\pm 0.018$
@@ 5.1 Location Finding
The location finding experiment was explored by Foster et al. (2021) and Blau et al. (2022). It is known there are $2$ objects placed in a 2-dimensional space, and the aim is to determine the unknown locations of these objects; denote the locations $\boldsymbol{\theta}=\{\boldsymbol{\beta}_{i}\}_{i=1}^{K}$, for $K=2$. Each object emits a signal which obeys an inverse-square law. We need to select designs $\xi$, which are coordinates in the space, in an effort to find the locations of the objects based on the observed signals. The signal strength of a single object increases as we select $\xi$ closer to this object, and it decays as we choose $\xi$ farther away. As we have multiple objects, we instead observe the total signal intensity, which is a superposition of the signals emitted by the individual objects. Hyperparameter details can be found in Appendix C, and experiment details can be found in Appendix D.1.
@@
SUNRISE-DroQ $13.938\pm 0.021$ $12.544\pm 0.022$ $21.210\pm 0.165$ $13.875\pm 0.049$ 18.16h
Random $11.145\pm 0.035$ $9.692\pm 0.030$ $12.047\pm 0.056$ $9.796\pm 0.032$
@@
Figure 2: Trained REDQ agent selecting designs at deployment time in the location finding experiment
@@ 5.2 Constant Elasticity of Substitution
The constant elasticity of substitution (CES) experiment was explored by Foster et al. (2019), Foster et al. (2020) and Blau et al. (2022). We have two baskets $\boldsymbol{x},\boldsymbol{x^{\prime}}\in[0,100]^{3}$ of goods, and a human indicates their preference of the two baskets on a sliding 0-1 scale. The CES model (Arrow et al. 1961) with latent variables $\boldsymbol{\theta}=(\rho,\boldsymbol{\alpha},u)$, which all characterise the human’s utility for the different items, is then used to measure the difference in utility of the baskets. The goal is to design the baskets in a way that allows for inference of the latent variables. The baskets are 3-tuples, meaning that we have $6$ design space dimensions $(\xi=(\boldsymbol{x},\boldsymbol{x^{\prime}}))$. Hyperparameter details can be found in Appendix C, and experiment details can be found in Appendix D.2.
@@ A.2 HiP-MDP
[…]
@@ B.2 Randomised Ensembled Double Q-Learning
REDQ extends SAC through employing various adjustments. This includes a higher UTD ratio (by allowing for $G$ updates as in the algorithm below), the use of an ensemble of Q-functions, and in-target minimisation over a random subset $M$ of the $N$ Q-functions (Chen et al. 2021).
Here, we can use $M$ Q-functions when computing the Q-target, which makes use of our ensemble of $N$ Q-functions. We could set $M=N$ as in SAC, although here we have the advantage of a large ensemble of Q-functions instead of always having 2 Q-functions. Using a small value of $M$, such as $M=2$, has been found to work well (Chen et al. 2021). Following from the SAC algorithm presented above, our Q-target is computed by,
@@
1:  Initialise policy parameters $\phi$, $N$ Q-function parameters $\theta_{i}$, and empty replay buffer $D$. Set target parameters $\theta_{\text{targ},i}\leftarrow\theta_{i}$. All for $i=1,\ldots,N$.
@@
8:        Compute the Q-target $y$ (same for all of the $N$ Q-functions) by
@@ B.3 Dropout Q-Functions for Doubly Efficient Reinforcement Learning
1:  Initialise policy parameters $\phi$, $N$ Q-function parameters $\theta_{i}$ and empty replay buffer $D$. Set target parameters $\theta_{\text{targ},i}\leftarrow\theta_{i}$. All for $i=1,\ldots,N$.
@@ B.4 Scaled-by-Resetting
1:  Initialise policy parameters $\phi$, $N$ Q-function parameters $\theta_{i}$, and empty replay buffer $D$. Set target parameters $\theta_{\text{targ},i}\leftarrow\theta_{i}$. All for $i=1,\ldots,N$. Set fixed reset interval $\psi$.
@@
8:        Compute the Q-target $y$ (same for all of the $N$ Q-functions) by
@@ B.5 Simple Unified Framework for Reinforcement Learning Using Ensembles
We explain the UCB exploration strategy (Chen et al. 2017; Lee et al. 2021) in full, which is an important component of our policy ensemble-based algorithm. The UCB exploration formula used in the SUNRISE framework (Lee et al. 2021) is designed to balance exploration and exploitation through using the $N$ candidate actions generated by our $N$ policies. The UCB formula is as follows,
@@
Consider an ensemble of $N$ SAC agents, $\{Q_{\theta_{i}},\pi_{\phi_{i}}\}_{i=1}^{N}$, where $\theta_{i}$ and $\phi_{i}$ denote the parameters of the $i$-th soft Q-function and policy, respectively. Conventional Q-learning is based on the Bellman backup, and so it can be affected by error propagation. This essentially means that errors in the previous Q-function induce “noise” to the true Q-value of the current Q-function. Therefore, for each agent $i$, a weighted Bellman backup (Lee et al. 2021) is used instead of $J_{Q}$ from SAC (Haarnoja et al. 2019), given by,
@@
1:  Initialise $N$ policy parameters $\phi_{i}$, $N$ Q-function parameters $\theta_{i}$, and empty replay buffer $D$. Set target parameters $\theta_{\text{targ},i}\leftarrow\theta_{i}$ and initial temperature parameters $\alpha_{i}$. All for $i=1,\ldots,N$.
@@
4:        Collect $N$ action samples: $A_{t}=\{a_{t,i}\sim\pi_{\phi_{i}}(a\mid s_{t})\mid i\in\{1,\ldots,N\}\}$
@@ C.2 REDQ Hyperparameters
REDQ $6.279\pm 0.013$ $11.689\pm 0.012$ $11.881\pm 0.013$ $11.507\pm 0.015$ $11.095\pm 0.017$ 13.23h
REDQ-Disc-0.99 $6.328\pm 0.012$ $11.536\pm 0.013$ $11.685\pm 0.014$ $11.250\pm 0.015$ $10.825\pm 0.017$ 13.41h
REDQ-Ens-10 6.423 $\pm$ 0.013 11.744 $\pm$ 0.013 $11.884\pm 0.013$ $11.419\pm 0.015$ $10.989\pm 0.017$ 38.91h
REDQ-Tau-0.01 $6.286\pm 0.013$ $11.651\pm 0.013$ 11.919 $\pm$ 0.013 11.516 $\pm$ 0.015 11.177 $\pm$ 0.016 13.38h
@@
REDQ-Disc-0.99 $6.331\pm 0.012$ $12.071\pm 0.018$ $12.704\pm 0.025$ $12.170\pm 0.027$ $11.653\pm 0.028$
REDQ-Ens-10 6.425 $\pm$ 0.013 12.427 $\pm$ 0.019 $13.140\pm 0.028$ $12.513\pm 0.029$ $11.943\pm 0.029$
@@ D.1 Location Finding
$\mu(\boldsymbol{\theta},\xi)=b+\sum_{i=1}^{K}\frac{\alpha}{m+||\boldsymbol{\beta}_{i}-\xi||^{2}},$
@@
where $\boldsymbol{0}$ is the mean vector, and $I$ is the covariance matrix, an identity matrix, both with dimension $d$.

### v2

@@ 3.2 Hidden-Parameter Markov Decision Process
  • The prior distribution, $P_{\boldsymbol{\Theta}}=p(\boldsymbol{\theta})$, of the model parameters. For our sequential BOED problems, $\boldsymbol{\theta}_{0:L}\overset{\text{i.i.d.}}{\sim}p(\boldsymbol{\theta})$;
@@
    where $C_{t}\cdot\boldsymbol{1}$ is a dot product with a vector of “11”s;
@@
To find the optimal policy we utilise deep RL algorithms [^3] In the experiments we conduct using these algorithms, we assume $\gamma$ values close to 11 (rather than $\gamma=1$) for computational stability; this follows Blau et al. (2022). (François-Lavet et al. 2018). The policy is implemented as a deep neural network that relies on a summary of $h_{t}$ in the form of an encoder network. Although the definition of the HiP-MDP satisfies the Markov property, it does so by explicitly including $h_{t}$ as part of $s_{t}$; the use of an encoder network weakens this by not explicitly requiring $h_{t}$ to be part of the state. Instead, the encoder network is recursively defined, so that a summary representation of $h_{t}$ at time $t$ only depends on the summary representation at time t1t-1 and $(\xi_{t},y_{t})$ (Blau et al. 2022). A detailed description of the policy network, as well as more details on the HiP-MDP implementation, are given in Appendix A.
@@ 4 Algorithms
REDQ $6.279\pm 0.013$ $1.689\pm 0.012 11$$881\pm 0.013 11.$$507\pm 0.015 11.$$1.095\pm 0.017$ 13.23h
SBR $6.165\pm 0.012$ $1.362\pm 0.013 11$$788\pm 0.014 11.$$518\pm 0.016 11.$$1.257\pm 0.017$ 13.52h
DroQ $6.368\pm 0.013$ $1.680\pm 0.013 11$$902\pm 0.013 11.$$480\pm 0.015 11.$$1.090\pm 0.017$ 19.05h
SUNRISE $6.340\pm 0.013$ 11.837 $\pm$ 0.012 $2.133\pm 0.013 11$.846 $\pm$ 0.014 $11.445\pm 0.016$ 21.44h
SUNRISE-DroQ 6.433 $\pm$ 0.012 $11.770\pm 0.012$ 12.143 $\pm$ 0.013 $1.831\pm 0.014 11$.453 $\pm$ 0.016 29.77h
Random $5.057\pm 0.010$ $7.443\pm 0.016$ $8.820\pm 0.018$ $9.730\pm 0.019$ $10.346\pm 0.018$ -
@@ 5.1 Location Finding
The location finding experiment was explored by Foster et al. (2021) and Blau et al. (2022). It is known there are 22 objects placed in a 2-dimensional space, and the aim is to determine the unknown locations of these objects; denote the locations $\boldsymbol{\theta}=\{\boldsymbol{\beta}_{i}\}_{i=1}^{K}$, for $K=2$. Each object emits a signal which obeys an inverse-square law. We need to select designs $\xi$, which are coordinates in the space, in an effort to find the locations of the objects based on the observed signals. The signal strength of a single object increases as we select $\xi$ closer to this object, and it decays as we choose $\xi$ farther away. As we have multiple objects, we instead observe the total signal intensity, which is a superposition of the signals emitted by the individual objects. Hyperparameter details can be found in Appendix C, and experiment details can be found in Appendix D.1.
@@
SUNRISE-DroQ $13.938\pm 0.021$ $544\pm 0.022 21.21$$210\pm 0.165$ $13.875\pm 0.049$ 18.16h
Random $11.145\pm 0.035$ $9.692\pm 0.030$ $12.047\pm 0.056$ $9.796\pm 0.032$ -
@@
Figure 2: Trained REDQ agent selecting designs at deployment time in the location finding experiment.
@@ 5.2 Constant Elasticity of Substitution
The constant elasticity of substitution (CES) experiment was explored by Foster et al. (2019), Foster et al. (2020) and Blau et al. (2022). We have two baskets $\boldsymbol{x},\boldsymbol{x^{\prime}}\in[0,100]^{3}$ of goods, and a human indicates their preference of the two baskets on a sliding 0-1 scale. The CES model (Arrow et al. 1961) with latent variables $\boldsymbol{\theta}=(\rho,\boldsymbol{\alpha},u)$, which all characterise the human’s utility for the different items, is then used to measure the difference in utility of the baskets. The goal is to design the baskets in a way that allows for inference of the latent variables. The baskets are 3-tuples, meaning that we have 66 design space dimensions $(\xi=(\boldsymbol{x},\boldsymbol{x^{\prime}}))$. Hyperparameter details can be found in Appendix C, and experiment details can be found in Appendix D.2.
@@ A.2 HiP-MDP
$B_{\psi,t}=\sum_{k=1}^{t}\text{Encoder}_{\psi}(\xi_{k},y_{k}),\hskip 4.30554ptC_{t}=\left[\prod_{k=1}^{t}p(y_{k}\mid\boldsymbol{\theta}_{\ell},\xi_{k})\right]^{L}_{\ell=0},$
@@ B.2 Randomised Ensembled Double Q-Learning
REDQ extends SAC through employing various adjustments. This includes a higher UTD ratio (by allowing for $G$ updates as in the algorithm below), the use of an ensemble of Q-functions, and in-target minimisation over a random subset $M$ of the NN Q-functions (Chen et al. 2021).
Here, we can use $M$ Q-functions when computing the Q-target, which makes use of our ensemble of NN Q-functions. We could set $M=N$ as in SAC, although here we have the advantage of a large ensemble of Q-functions instead of always having 2 Q-functions. Using a small value of $M$, such as $M=2$, has been found to work well (Chen et al. 2021). Following from the SAC algorithm presented above, our Q-target is computed by,
@@
1: Initialise policy parameters $\phi$, NN Q-function parameters $\theta_{i}$, and empty replay buffer $D$. Set target parameters $\theta_{\text{targ},i}\leftarrow\theta_{i}$. All for $i=1,\ldots,N$.
@@
8:   Compute the Q-target $y$ (same for all of the NN Q-functions) by
@@ B.3 Dropout Q-Functions for Doubly Efficient Reinforcement Learning
1: Initialise policy parameters $\phi$, NN Q-function parameters $\theta_{i}$ and empty replay buffer $D$. Set target parameters $\theta_{\text{targ},i}\leftarrow\theta_{i}$. All for $i=1,\ldots,N$.
@@ B.4 Scaled-by-Resetting
1: Initialise policy parameters $\phi$, NN Q-function parameters $\theta_{i}$, and empty replay buffer $D$. Set target parameters $\theta_{\text{targ},i}\leftarrow\theta_{i}$. All for $i=1,\ldots,N$. Set fixed reset interval $\psi$.
@@
8:   Compute the Q-target $y$ (same for all of the NN Q-functions) by
@@ B.5 Simple Unified Framework for Reinforcement Learning Using Ensembles
We explain the UCB exploration strategy (Chen et al. 2017; Lee et al. 2021) in full, which is an important component of our policy ensemble-based algorithm. The UCB exploration formula used in the SUNRISE framework (Lee et al. 2021) is designed to balance exploration and exploitation through using the NN candidate actions generated by our NN policies. The UCB formula is as follows,
@@
Consider an ensemble of NN SAC agents, $\{Q_{\theta_{i}},\pi_{\phi_{i}}\}_{i=1}^{N}$, where $\theta_{i}$ and $\phi_{i}$ denote the parameters of the $i$-th soft Q-function and policy, respectively. Conventional Q-learning is based on the Bellman backup, and so it can be affected by error propagation. This essentially means that errors in the previous Q-function induce “noise” to the true Q-value of the current Q-function. Therefore, for each agent $i$, a weighted Bellman backup (Lee et al. 2021) is used instead of $J_{Q}$ from SAC (Haarnoja et al. 2019), given by,
@@
1: Initialise NN policy parameters $\phi_{i}$, NN Q-function parameters $\theta_{i}$, and empty replay buffer $D$. Set target parameters $\theta_{\text{targ},i}\leftarrow\theta_{i}$ and initial temperature parameters $\alpha_{i}$. All for $i=1,\ldots,N$.
@@
4:   Collect NN action samples: $A_{t}=\{a_{t,i}\sim\pi_{\phi_{i}}(a\mid s_{t})\mid i\in\{1,\ldots,N\}\}$
@@ C.2 REDQ Hyperparameters
REDQ $6.279\pm 0.013$ $1.689\pm 0.012 11$$881\pm 0.013 11.$$507\pm 0.015 11.$$1.095\pm 0.017$ 13.23h
REDQ-Disc-0.99 $6.328\pm 0.012$ $1.536\pm 0.013 11$$685\pm 0.014 11.$$1.250\pm 0.015$ $10.825\pm 0.017$ 13.41h
REDQ-Ens-10 6.423 $\pm$ 0.013 11.744 $\pm$ 0.013 $1.884\pm 0.013 11$$1.419\pm 0.015$ $10.989\pm 0.017$ 38.91h
REDQ-Tau-0.01 $6.286\pm 0.013$ $1.651\pm 0.013 11$.919 $\pm$ 0.013 11.516 $\pm$ 0.015 11.177 $\pm$ 0.016 13.38h
@@
REDQ-Disc-0.99 $6.331\pm 0.012$ $12.071\pm 0.018$ $12.704\pm 0.025$ $2.170\pm 0.027 11$$1.653\pm 0.028$
REDQ-Ens-10 6.425 $\pm$ 0.013 12.427 $\pm$ 0.019 $13.140\pm 0.028$ $2.513\pm 0.029 11$$1.943\pm 0.029$
@@ D.1 Location Finding
[…]
@@
where $\boldsymbol{0}$ is the mean vector, and II is the covariance matrix, an identity matrix, both with dimension $d$.